Hayden has two coin banks containing only nickles and quarters. The total value of the coins in the first bank is $2.90. In the second bank, Hayden has two more quarters than in the first bank and twice as many nickles. the value of the coins in the second bank is $3.80. Determine the number of nickles and quarters Hayden has in the second bank.
Answer by lwsshak3(11628)

Hayden has two coin banks containing only nickles and quarters. The total value of the coins in the first bank is $2.90. In the second bank, Hayden has two more quarters than in the first bank and twice as many nickles. the value of the coins in the second bank is $3.80. Determine the number of nickles and quarters Hayden has in the second bank.
***
let x=number of nickels in the first bank
let y=number of quarters in the first bank
2x=number of nickels in the second bank
y+2=number of quarters in the second bank
..
.05x+.25y=2.90 (first bank)
.05(2x)+.25(y+2)=3.80 (second bank)
..
.05x+.25y=2.90
.10x+.25y+.50=3.80
..
.05x+.25y=2.90
.10x+.25y=3.30
subtract:
-.05x=-.40
x=8
.25y=2.90-.05x=2.90-.40=2.50
y=10
..
2x=16
y+2=12
number of nickels in the second bank=16
number of quarters in the second bank=12